OpenStudy (anonymous):

a ball is dropped from rest from a tower and strikes the ground 125m below. Approxmiately how many seconds does it take for the ball to strike the ground after being dropped. Neglect air resistance

OpenStudy (anonymous):

gravity is being used in this one, right?

OpenStudy (anonymous):

so change a with g

OpenStudy (anonymous):

let's start given: Vi=0 m/s acceleration due to gravity= 9.8 m/s^2 displacement: 125m time=?

OpenStudy (anonymous):

yes, you're right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=0 right?

OpenStudy (anonymous):

and y initial=125?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so which formula are we going to use? we don't have \(V_f\) so...?

OpenStudy (anonymous):

I got -4.9t squared+t+125=0

OpenStudy (anonymous):

so does this mean time for quadtatrics?

OpenStudy (anonymous):

or did I mess up?

OpenStudy (anonymous):

I got 5.15 s

OpenStudy (anonymous):

is relatively close to 5.05 s, the answer

OpenStudy (anonymous):

i want to do it, step by step , so bare with me XD so i will use \[\Delta d=V _{i}\Delta t + \frac{ 1 }{2 } a \Delta t ^2\] Vi=0 so it will cancel out it will become \[\Delta d=\frac{ 1 }{2 } a \Delta t ^2\] then evaluate, to solve for \(\Delta t\) so it will be (\Delta t=\sqrt {\frac{2 \Delta d}{a}}\)

OpenStudy (anonymous):

\(\LARGE \Delta t=\sqrt {\frac{2 \Delta d}{a}}\)

OpenStudy (anonymous):

did you get what i did?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

wat was your final answer?

OpenStudy (anonymous):

I got time by using the quadratic formula

OpenStudy (anonymous):

which was 5.05s, the answer

OpenStudy (anonymous):

approx. 5.05 s

OpenStudy (anonymous):

YEAH, my hero!!!!!

OpenStudy (anonymous):

lowl, i'm practicing too XD nice questions

OpenStudy (anonymous):

ok, next problem?

OpenStudy (anonymous):

ok! ^.^